A particular fruit's weights are normally distributed, with a mean of 790 grams and a standard deviation of 21 grams.

If you pick 14 fruit at random, what is the probability that their mean weight will be between 773 grams and 794 grams
Answer #1

Solution :

\sigmaT = \sigma / \sqrt n = 21 / \sqrt 14 = 5.6125

= P[(773 - 790) / 5.6125 < (\bar x - \mu T) / \sigma T < (794 - 790) / 5.6125)]

= P(-3.0290 < Z < 0.7127)

= P(Z < 0.7127) - P(Z <-3.0290 )

= 0.762 - 0.0012

= 0.7608  

Probability = 0.7608